OpenStudy (anonymous):

PLEASE HELP ME ****need this to be 100% right please***

OpenStudy (anonymous):

@Nurali

OpenStudy (anx):

Pentagon

OpenStudy (uri):

which test?

OpenStudy (anonymous):

thank you @anx

OpenStudy (anx):

heres the proof 180(n-2) pentagon has 5 sides so n = 5 180(5-2) = 180(3)= 540

OpenStudy (anonymous):

oh thank you

OpenStudy (anx):

yup